You must show your work to get any credit.

Anderson Corp just issued a stock dividend of$1.00 yesterday. The company plans on increasing the dividend by 5% per year for the next 5 years. After which, the dividend will grow at 4% forever. The required rate of return is 8%.

A) Calculate the current price of the stock

B) Calculate what the price of the stock should be in 1 year

C) Calculate the Dividend yield and capital gains yield during the first year.
Answer #1

a.

Stock Price = PV(Dividend) + PV(Horizon Value)

Stock Price = 1(1.05/1.08) + 1(1.05/1.08)2 + 1(1.05/1.08)3 + 1(1.05/1.08)4 + 1(1.05/1.08)5 + 1(1.05/1.08)5(1.04)/(0.08 - 0.04)

Stock Price = 0.97 + 0.95 + 0.92 + 0.89 + 0.87 + 22.58

Stock Price = $27.18

b.

Stock Price in Year 1 = 1(1.05/1.08) + 1(1.05/1.08)2 + 1(1.05/1.08)3 + 1(1.05/1.08)4 + 1(1.05/1.08)4(1.04)/(0.08 - 0.04)

Stock Price in Year 1 = 0.97 + 0.95 + 0.92 + 0.89 + 23.23

Stock Price in Year 1 = $26.96